OpenStudy (anonymous):

Use mathematical induction to prove the statement is true for all positive integers n, or show why it is false. 1^2 + 4^2 + 7^2+...+ (3n-2)^2 = n(6n^2-3n-1)/2

OpenStudy (anonymous):

@phi

OpenStudy (phi):

First, prove the base case. Is the formula true for n=1 ?

OpenStudy (anonymous):

Yes

OpenStudy (phi):

so write that part down. base case is true. now assume the formula is correct for n= k is it true for k+1 ? to do this part involves algebra

OpenStudy (anonymous):

Yes, that's where I get stuck. I don't know how to do that.

OpenStudy (phi):

on the right side, replace n with (k+1)

OpenStudy (anonymous):

For proving for n=k+1 what you do is this. You've assumed n=k is correct so write what you got for assuming n=k down. Then add on 1 more term of (3n-2)^2 where n=k+1.

OpenStudy (anonymous):

So would the new equation be (k+1)(6(k+1)^2-3(k+1)-1)/2?

OpenStudy (phi):

on the left side, show the last "valid" term (replace n with k) and add one more term, with n replaced with k+1

OpenStudy (anonymous):

No. You need to put in the equation when n=k so k(6k^2-3k-1)/2 then add on one more term where n=k+1 so 3((k+1)-2)^2.

OpenStudy (anonymous):

I am so lost.

OpenStudy (anonymous):

Right, when you assume n=k you get.\[\sum_{r=1}^{k}(3r-2)^2 = k(6k^2 - 3k - 1)/2\]

OpenStudy (anonymous):

For proving n=k+1 you put that Right hand equation back in and just add on one more term of (3r-2)^2 where r=k+1

OpenStudy (anonymous):

So it would be \[k(6k^2 - 3k - 1)/2 + (3k+3-2)^2\]

OpenStudy (anonymous):

And you want to prove that that is equal to n(6n^2-3n-1)/2 when n=k+1

OpenStudy (anonymous):

(k+1)(6(k+1)^2-3(k+1)-1)/2. You're trying to prove it's equal to that

OpenStudy (anonymous):

Okay, would I replace k with 1?

OpenStudy (anonymous):

No. You need to expand the brackets and then factorize out the correct factors which will be k+1 and 1/2

OpenStudy (anonymous):

How do I do that?

OpenStudy (anonymous):

\[\frac{ 6k^3 - 3k^2 - k }{ 2 } + (3k +1)^2\]

OpenStudy (anonymous):

Okay, then what?

OpenStudy (anonymous):

Expand the (3k+1)^2

OpenStudy (anonymous):

Do I FOIL (3k+1)(3k+1)?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

9k^2 + 6k +2

OpenStudy (phi):

+1 not +2

OpenStudy (anonymous):

Ah, okay. Then what?

OpenStudy (anonymous):

Factorize the 1/2

OpenStudy (anonymous):

English please? Lol

OpenStudy (anonymous):

take 1/2 out as a factor

OpenStudy (anonymous):

Can I ask why you're doing induction?

OpenStudy (anonymous):

Because I'm in pre-calculus...

OpenStudy (anonymous):

But, I got 1/2(18k^2+12k+4)

OpenStudy (anonymous):

+2

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

And then add in the first half that you've factorized also

OpenStudy (anonymous):

1/2(6k^3-3k^2-k+18k^2+12k+2)

OpenStudy (anonymous):

Then collect the terms.

OpenStudy (anonymous):

1/2(6k^3+15k^2+11k+2)

OpenStudy (anonymous):

Then factorize out k+1

OpenStudy (phi):

now simplify the right side and show it equals that expression

OpenStudy (anonymous):

This is a real b*tch to simplify which is annoying.

OpenStudy (anonymous):

If I were you, I'd get someone who knows how to do proof by induction to go through it with you in person on paper because it is much easier like that than trying to explain with type.

OpenStudy (phi):

(6k^3+15k^2+11k+2)/2 = ? (k+1)(6(k+1)^2-3(k+1)-1)/2 or 6k^3+15k^2+11k+2 = ? (k+1)(6(k+1)^2-3(k+1)-1) you are trying to show the left side equals the right side. perhaps the easiest way is to expand the right side, collect terms...

OpenStudy (phi):

The left side came from the work you and alright just did the right side came from putting k+1 into the formula working on the right side (k+1)(6(k+1)^2-3(k+1)-1) concentrate on (k+1)^2 that is (by foil) k^2+2k+1 so we have (k+1)( 6( k^2+2k+1) - 3(k+1) -1) can you finish ?